OpenStudy (anonymous):

Rad(1/3)Rad(-27)... answer is 3i.. how?

OpenStudy (anonymous):

I would combine everything into one radical

OpenStudy (anonymous):

take the -1 out of the radical

OpenStudy (anonymous):

rad(-1) is i btw

OpenStudy (anonymous):

\[\sqrt{\frac{1}{3}} \sqrt{-27} = \sqrt{-27 * \frac{1}{3}} = \sqrt{-9} = \sqrt{-1}\sqrt{9}=isqrt{9}=i*3=3*i\]